OpenStudy (anonymous):

Find the product of (x + 5)2. x2 − 10x + 25 x2 − 25 x2 + 25 x2 + 10x + 25

OpenStudy (anonymous):

2x+10

OpenStudy (anonymous):

explain.

OpenStudy (skullpatrol):

You should type a "^" if you mean $$(x+5)^2$$

OpenStudy (skullpatrol):

(x + 5)2 = (x+5)*2

OpenStudy (anonymous):

ok

OpenStudy (skullpatrol):

@mikaa_toxica13 As long as you understand; that is what is important ;-)

OpenStudy (anonymous):

okay well i never really said i did :/

OpenStudy (skullpatrol):

Please ask questions :-(

OpenStudy (anonymous):

okay, can u please fully explain it

OpenStudy (skullpatrol):

I did explain the FOIL method to you remember?

OpenStudy (anonymous):

yes

OpenStudy (skullpatrol):

These type of questions are a "special case" of that general method.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

can u explain this particular question please.

OpenStudy (skullpatrol):

Sure.

OpenStudy (skullpatrol):

$$(x + 5)^2 = (x+5)(x+5)$$

OpenStudy (skullpatrol):